Quantitative Problem: Rosnan Industries 2019 and 2018 balance sheets and income statements are shown below. Balance Sheets:Shares outstanding Price WACC 100 $25.00 10.00 % 100 $22.50 What is the firms 2019 current ratio? Round your answer to two d
Answer #1

what is the firm's 2019 current ratio?

current ratio = current assets/ current liabilities

current asset =$740

Current Liabilities =$365

Current ratio= $740 /$365 = 2.03

The 2019 current ratio indicates that Rosnan has sufficient current assets to meet its current obligations as they come due.

what is the firm's 2019 total assets turnover ratio?

Total Assets Turnover ratio = Sales / Total assets

Sales=$1,885

Total assets =$2,740

Total Asset Turnover = $1,885/$2,740 = 0.6879

current asset turnover =Sales / Current Assets = $1,885/$740 =2.5472

Fixed asset turnover = Sales / fixed assets = $1,885/$2,000=0.9425

Given the 2019 current and total assets turnover ratios calculates above, if Rosnan's 2019 quick ratio is 1.0 then an analyst might conclude that Rosnan's fixed assets are managed efficiently

what is the firm's 2019 debt to capital ratio?

Debt to capital = total liabilities/ (total liabilities + equity)

Debt to capital = $365+450/$2,740=$815/$2,740 = 0.2974 or 29.74%

If the industry average debt to capital is 30%, then Rosnan's creditors have more cushion than indicated by the industry average.

what is the firm's Profit margin?

Profit margin = Net Income / sales * 100

Net income =$354

Sales =$ 1,885

Profit Margin =$354/$1,885*100 = 18.77%

if the industry average profit margin is 12%,then Rosnan's lower than average debt to capital ratio might be one reason for its high profit margin

what is the firm's 2019 price/ earning ratio?

price / earning ratio = share price/ EPS

EPS = Net income / No. of share = $354/100 = $3.54

Price / earning ratio = $25/$3.54 = $7.06

Using the DuPont equation, what is the firm's 2019 ROE?

ROE = Profit margin * Total assets turnover * equity multiplier

Profit Margin = 18.77%

Total Asset Turnover = 0.687

Equity Multiplier= 1/ (1-debt to capital ratio)

=1/ (1-0.297)= 1/ 0.703 = 1.422

ROE = 18.77%* 0.687*1.422 = 0.1833 or 18.33%